Group Selection Page
HELP
Controls Group
Data Analysis Group
ITDM
WP76
DET
Vacuum
EEE
WP78
WP74
FXE
SCS
SPB
MID
HED
SQS
SXP
Sample Environment
Operation
PSPO
XO
TS
Migrated2Zulip
|
Technical integration and coordination
SASE1 Status and SASE3 Readiness
|
Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther |
Not logged in |
 |
|
Message ID: 569
Entry time: 27 Nov 2024, 09:44
|
Author: |
Frederik Wolff-Fabris |
Location: |
General |
Severity: |
Information |
Subject: |
Notes XO Integration Meeting 27.11.2024 |
|
|
General Announcement:
- Issues with machine at the cold compressor happened and the restart is delayed. Expectation for the cryo stable is around 15:00 today.
Tuning will start from this point and non-standard modes probably be tuned tomorrow. Further info will be distributed during the day.
CTRL:
- Nothing to report.
DA:
- Nothing to report.
EEE:
- Yesterday a ZZ access related to the SA1 EPS issues; power supply was exchanged and filters and motors are operational again. Tests on the broken components will be done.
- L. Zanellato adds that cables shall not be the origin of the problem.
VAC:
- Nothing to report.
XPD:
- YAG screen for the SA1 FEL imager was placed in a new spot and is not anymore visible.
DET:
- LPD gain: follow up with CTRL is to be done.
ITDM:
- Nothing to report.
XRO:
- Nothing to report.
LAS:
- Nothing to report.
UNSYS:
- Nothing to report.
SPB:
- Nothing special since Monday.
FXE:
- Nothing to report.
- Difficulties in using the beam over the weekend for extended hours: no further information.
MID:
- Nothing to report.
HED:
- Nothing to report.
SCS:
- Nothing to add.
SQS:
- Nothing to report.
SXP:
- Nothing to report. |